OpenStudy (deadshot):

What is the simplified form of x plus 2 over x-squared minus 3x minus 10 • x minus 3 over x-squared plus x minus 12 ?

OpenStudy (anonymous):

\[\frac{ x+2 }{x ^{2}-3x-10 }.\frac{ x-3 }{ x ^{2}+x-12}\] \[\frac{ x+2 }{ x ^{2}-5x+2x-10 }.\frac{ x-3 }{x ^{2} +4x-3x-12}\] \[\frac{ x+2 }{x(x-5)+2(x-5)}.\frac{ x-3 }{ x(x+4)-3(x+4) }\] \[\frac{ x+2 }{ (x-5)(x+2) }.\frac{ x-3 }{ (x+4)(x-3)}\]

OpenStudy (anonymous):

now this is the simplified form: \[\frac{ 1 }{ (x-5)(x+4) }\]

OpenStudy (deadshot):

Thanks!